TIME-RESOLVED RECOMBINATION DYNAMICS OF IBr^{−}(CO_{2})_{n} CLUSTERS

\underline{J.P. MARTIN}, V. DRIBINSKI, J. BARBERA, M.A. THOMPSON, R. PARSON, and W.C. LINEBERGER, JILA/Department of Chemistry and Biochemistry, University of Colorado, Boulder, CO 80309; A.SVENDSEN, Department of Physics and Astronomy, University of Aarhus, Aarhus, Denmark.

We report the femtosecond dynamics of IBr^{−}(CO_{2})_{n} (n = 5, 6, 7, 8, 10) clusters. Excitation to the dissociative IBr^{−} A′ 2Π_{1}_{/}_{2} state of the chromophore is achieved with a 180 fs 795 nm laser pulse. Dissociation from the A′ state of the bare anion results in I^{−} and Br products. Solvent molecules surrounding the dihalide chromophore promote recombination of the dissociating IBr^{−}, in part through a solvent-induced well on the A′ state that can trap the evolving cluster. The recombination time is determined by using a delayed femtosecond probe laser at the same wavelength, detecting recombined IBr^{−} cluster ions. In sharp contrast to previous I_{2}^{−} studies, the observed recombination times increase dramatically with increasing cluster size, from 15 ps for n = 5 to 900 ps for n = 10. Extensive electronic structure calculations and non-adiabatic molecular dynamics simulations provide a framework to understand this unexpected behavior. Supported by NSF and AFOSR